OpenStudy (anonymous):

2n+3=5n-5 solve for n

OpenStudy (anonymous):

move variables.. 8=3n n=8/3

OpenStudy (anonymous):

how did you move the variabless

OpenStudy (anonymous):

2n+3 = 5n-5 2n = 5n - 5 - 3 2n = 5n - 8 2n + 8 = 5n 8 = 5n - 2n 8 = 3n n = 3/8

OpenStudy (anonymous):

sorry n = 8/3